square root(x (square root x (square root x))) = ?
a) x^7/8 b) x^7/4 c) x^15/16 d) x^3/4 e) x^15/8

Answer by stanbon(75887)

square root(x (square root x (square root x))) = ?
a) x^7/8 b) x^7/4 c) x^15/16 d) x^3/4 e) x^15/8
----
= square root(x (square root x (x^(1/2))) = ?
---
= square root(x (square root x^(3/2))))
---
= sqrt(x (x^(3/4))
---
= sqrt(x^(7/4))
---
= x^(7/8)
=====================
Cheers,
Stan H.
Reply to stanbon@comcast.net